# Spark Transformer: Reactivating Sparsity in FFN and Attention

### v1

@@ Abstract
The discovery of the lazy neuron phenomenon [49] in trained Transformers, where the vast majority of neurons in their feed-forward networks (FFN) are inactive for each token, has spurred tremendous interests in activation sparsity for enhancing large model efficiency. While notable progress has been made in translating such sparsity to wall-time benefits across CPUs, GPUs, and TPUs, modern Transformers have moved away from the ReLU activation function crucial to this phenomenon. Existing efforts on re-introducing activation sparsity, e.g., by reverting to ReLU, applying top-$k$ masking, or introducing a sparse predictor, often degrade model quality, increase parameter count, complicate or slow down training. Sparse attention, the application of sparse activation to the attention mechanism, often face similar challenges.
This paper introduces the Spark Transformer, a novel architecture that achieves a high level of activation sparsity in both FFN and the attention mechanism while maintaining model quality, parameter count, and standard training procedures. Our method realizes sparsity via top-$k$ masking for explicit control over sparsity level. Crucially, we introduce statistical top-$k$, a hardware-accelerator-friendly, linear-time approximate algorithm that avoids costly sorting and mitigates significant training slowdown from standard top-$k$ operators. Furthermore, Spark Transformer reallocates existing FFN parameters and attention key embeddings to form a low-cost predictor for identifying activated entries. This design not only mitigates quality loss from enforced sparsity, but also enhances wall-time benefit. Pretrained with the Gemma-2 recipe, Spark Transformer demonstrates competitive performance on standard benchmarks while exhibiting significant sparsity: only 8% of FFN neurons are activated, and each token attends to a maximum of 256 tokens. This sparsity translates to a 2.5$\times$ reduction in FLOPs, leading to decoding wall-time speedups of up to 1.79$\times$ on CPU and 1.40$\times$ on GPU.
@@ 1 Introduction
The machine learning field has experienced a rapid expansion in the scale of Transformer models [4, 3, 24, 1], significantly advancing the state-of-the-art in language understanding and generation. However, this pursuit of larger models is often constrained not by inherent limitations in model quality [40], but by the soaring computational costs [71, 63] associated with the number of parameters. This challenge is compounded by the trend towards models capable of processing increasingly longer input sequences [67], where computational demands scale proportionally with input length.
@@
Figure 1: Spark Transformer improves inference efficiency via high-level activation sparsity in both FFN and attention, while being nearly quality neutral. (a) Comparison to prior work in terms of relative FLOPs per token at 8k sequence length (y-axis) vs relative training loss (x-axis) [^2] All models in Figure 1(a) were trained with standard Gemma-2 recipe for 1/6 of the full pretraining iterations. All other results in this paper are obtained for a fully pretrained Spark Transformer. . [$\blacksquare$] We use standard Gemma-2 [27] as baseline, which has no activation sparsity. [] Methods employing activation sparsity within the FFN layers only. Our Spark FFN achieves the most favorable trade-off compared to ReLU, ReLU2, and Topk, which refer to standard Gemma-2 with activation function switched to ReLU [59], ReLU2 [86], and the composition of Topk and GELU, respectively. [$\blacktriangle$] Combining Spark FFN (with 8% activated parameters) with Spark Attention (with at most $256$ attended tokens), our Spark Transformer achieves performance comparable to Gemma-2 while reducing FLOPs to 40%. (b) Evaluation on standard downstream tasks confirms near-quality neutrality of Spark Transformer. (c) Prefill / decode wall time demonstrate a 1.86$\times$/1.64$\times$ speedup resulting from FLOPs reduction. Results are obtained on a 4-Core CPUs for prompts of 4096 tokens. For prefill, the prompt is chunked into batches of 64 tokens, following a default setup of gemma.cpp [30].
Activation sparsity has emerged as a prominent technique for mitigating the computational burdens associated with both large model sizes and long input sequences. For large models, activation sparsity reduces computational cost by activating only a small fraction of the model’s parameters for each input. This approach has attracted considerable attention due to the observed lazy neuron phenomenon [49] where natural activation sparsity occurs in the feed-forward networks (FFNs) of traditional Transformer models like T5 [66] and ViT [23], without explicit enforcement [85]. Subsequent work has successfully demonstrated wall-time benefits from activation sparsity on multiple hardware platforms, including CPU [84], GPU [77, 53], and TPU [83].
Despite the great success, a fundamental challenge arises in the application of sparse activation to the latest and state-of-the-art models. That is, the inherent sparsity, which is pivotal for obtaining efficiency, is largely absent due to the adoption of gated non-ReLU activation functions [18], widely adopted in, e.g., Mistral [37], Gemma [27], LLAMA [24]. This motivates the following question of reactivating sparsity:
@@
To answer this question, prior work has explored reverting to ReLU and its variants [59, 86], incorporating top-$k$ thresholding [83, 76, 78], or introducing a low-cost sparsity predictor [53, 84, 77, 83]. However, (Challenge #1) these methods often lead to a degradation in model quality (see Figure 1(a)). Furthermore, (Challenge #2) top-$k$ thresholding based methods involve sorting to obtain maximally activated neurons, which is inefficient on ML accelerators, such as TPUs [39], with possibly a 10×10$\times$10 × training slowdown (see Figure 6). Finally, (Challenge #3) the inclusion of a sparsity predictor often increases training pipeline complexity and incurs additional training costs and parameters.
Sparse attention, the application of activation sparsity to the attention mechanism, presents similar challenges. A common strategy is top-$k$ attention [31], which applies a top-$k$ mask to the attention coefficients. This can be augmented with a low-cost predictor to further enhance efficiency [69, 82, 46]. Nevertheless, achieving both high sparsity and accurate prediction without resorting to complex procedures, training slowdown from top-$k$, and compromising quality remains an open problem.
@@
Specifically, the Spark Transformer comprises Spark FFN and Spark Attention, both of which exploit the interpretation of FFNs and attention mechanisms as key-value lookup tables [29] to provide a unified framework for sparsity and a low-cost predictor (see Figure 2). Our low-cost predictor is constructed by repurposing a subset of the dimensions of the query and key vectors to compute an importance score for each key-value pair (see Section 2). This design addresses Challenge #3 by avoiding the introduction of extra parameters, enabling all model parameters to be trained in a single stage. We then introduce a statistical top-$k$ operator, which is applied to these predicted scores to select the activated keys. In particular, statistical top-$k$ is a linear-complexity algorithm for approximate nearest neighbor search, which addresses the issue of high computation cost of standard top-$k$ algorithms (i.e., Challenge #2). As explained in Section 3, this is achieved by fitting a Gaussian distribution to the activation scores and estimating a threshold that selects the top entries.
Finally, we demonstrate that Spark Transformer alleviates quality loss (i.e., Challenge #1) by performing a full pretraining using the Gemma-2 recipe [27]. Comparison with prior work shows that Spark Transformer exhibits a more favorable trade-off between FLOPs reduction and quality measured by pretraining loss (Figure 1(a)). Further evaluation on standard benchmarks confirms that Spark Transformer closely matches the performance of Gemma-2 (see Figure 1(b)), despite exhibiting a high degree of sparsity: only 8% of FFN neurons are activated, and each token attends to at most 256 tokens. Leveraging this sparsity, we assess the model’s inference efficiency on CPUs, demonstrating wall-time speedups of 1.86×1.86$\times$1.86 × for prefill and 1.64×1.64$\times$1.64 × for decode (see Figure 1(c)). Furthermore, an up to 1.4×1.4$\times$1.4 × wall-time speedup is obtained on NVIDIA T4 GPU (see Section 4). This enhanced efficiency broadens access to high-quality models for users with limited access to high-FLOP hardware, such as high-end GPUs and TPUs.
@@ 2.1 Spark FFN
Each matrix multiplication in Equation 1 has $2d_{\text{model}}\cdot d_{\text{ff}}$ FLOPs hence overall the computation cost is $4d_{\text{model}}\cdot d_{\text{ff}}$. Previous work shows that when $\sigma()$ is ReLU, the activation map $\sigma({\bm{K}}^{\top}{\bm{q}})$ is very sparse after model training. The sparsity can be used trivially to reduce the computation costs in the calculation of its product with the second layer weight matrix ${\bm{V}}$ [49], reducing the overall FLOPs count of FFN to $2d_{\text{model}}\cdot(d_{\text{ff}}+k)$, where $k\ll d_{\text{ff}}$ is the number of nonzero entries in the activation. Note that the sparsity cannot be used to reduce the computation costs associated with ${\bm{K}}$, which constitute half of the total FLOPs in FFN.
Figure 2: Architecture of Spark FFN and Spark Attention. (Left) Unified illustration of standard FFN (i.e., Equation 1) and standard Attention (i.e., Equation 6). In the case of FFN, ${\bm{q}}\in\mathbb{R}^{d_{\text{model}}}$ is the input, ${\bm{K}}$ and ${\bm{V}}$ are the first and second layer weights, respectively, and $\sigma()$ is GELU. In the case of Attention, ${\bm{q}}\in\mathbb{R}^{d_{\text{attn}}}$ is the query, ${\bm{K}}$ and ${\bm{V}}$ are key and value matrices, respectively, and $\sigma()$ is softmax. (Right) Unified illustration of Spark FFN (i.e., Equation 2) and Spark Attention (i.e., Equation 7). In the case of Spark FFN, $\sigma_{1}()$ is GELU and $\sigma_{2}()$ is identity. In the case of Spark Attention, $\sigma_{1}()$ is softmax and $\sigma_{2}()$ is softplus. In both cases, $\operatorname{{Statistical-Top}}_{k}$ (i.e., Equation 10) is applied to introduce sparsity, which enables sparse matrix multiplication with ${\bm{K}}_{2}$ and ${\bm{V}}$ that reduces FLOPs count.
@@
In above, ${\bm{K}}_{1}\subseteq\mathbb{R}^{r\times d_{\text{ff}}}$, ${\bm{K}}_{2}\subseteq\mathbb{R}^{(d_{\text{model}}-r)\times d_{\text{ff}}}$, and ${\bm{V}}\subseteq\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$ are trainable parameters, and the activation $\sigma()$ is taken to be GELU [33] following Gemma. $\operatorname{Top}_{k}$ is introduced for obtaining sparsity, with $k$ being a hyper-parameter specifying the sparsity level. Concretely, $\operatorname{Top}_{k}$ preserves the largest $k$ values in the activation vector, while setting other values to 0. In this paper, we use the efficient statistical top-k presented in Section 3, which avoids sorting activation values. Finally, the input ${\bm{q}}$ is split into ${\bm{q}}[:\!r]$ and ${\bm{q}}[r\!:]$, which contain the first $r$ and the rest of the dimensions, respectively, with $r$ being a hyper-parameter. It is introduced so that the term ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ serves as a low-rank predictor of the location of the nonzero entries, which allows us to obtain efficiency benefits in computing ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ and the multiplication with ${\bm{V}}$. This is discussed in detail below.
FLOPs per Token. Direct implementation of the Spark-FFN without exploiting sparsity has the same number of FLOPs as the standard FFN in Equation 1, i.e.,
@@
where the three terms are from multiplication with ${\bm{K}}_{1}$, ${\bm{K}}_{2}$, and ${\bm{V}}$, respectively. In Spark-FFN, one may first compute the term ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ as a low-rank predictor. After passing its output through $\operatorname{Top}_{k}$, which selects approximately the $k$ most important entries, followed by the activation function $\sigma()$, we obtain a sparse output. Importantly, after obtaining the sparse output there is no need to perform the full computation of the other two matrix multiplications in Equation 2, i.e., ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ and the multiplication with ${\bm{V}}$. Instead, one can perform a sparse matrix multiplication with a drastically reduced FLOPs count:
@@
which is an increasing function of $r$. In other words, $r$ controls the computation cost. We provide ablation study in Section C.4 to show that the best model quality is obtained when $r\approx\frac{d_{\text{model}}}{2}$. In this case, the total FLOP count of Spark FFN is approximately $d_{\text{model}}\cdot d_{\text{ff}}+3\cdot d_{\text{model}}\cdot k$, which is a $4$-times reduction from Equation 3 when $k$ is small.
Relation to gated activation. Many of the most recent Transformers, including Gemma, use a variant of the standard FFN in Equation 1 where the activation function is replaced with a gated one:
@@
In above, $\{{\bm{K}}_{1},{\bm{K}}_{2},{\bm{V}}\}\subseteq\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}^{\prime}}$. Note that when compared with the FFN in Equation 1 for quality studies, $d_{\text{ff}}^{\prime}$ is usually taken to be $2/3\cdot d_{\text{ff}}$ to be iso-parameter count [72].
Our Spark FFN in Equation 2 bears some resemblance to Gated FFN in that both have two linear maps in the first layer and one in the second layer. The difference lies in that 1) Spark FFN adds a $\operatorname{Top}_{k}$ to obtain sparsity, and 2) the input to the first layers of Spark FFN are obtained from splitting the dimensions of the input.

### v2

@@ Abstract
The discovery of the lazy neuron phenomenon Li et al. (2022) in trained Transformers, where the vast majority of neurons in their feed-forward networks (FFN) are inactive for each token, has spurred tremendous interests in activation sparsity for enhancing large model efficiency. While notable progress has been made in translating such sparsity to wall-time benefits across CPUs, GPUs, and TPUs, modern Transformers have moved away from the ReLU activation function crucial to this phenomenon. Existing efforts on re-introducing activation sparsity, e.g., by reverting to ReLU, applying top-$k$ masking, or introducing a sparse predictor, often degrade model quality, increase parameter count, complicate or slow down training. Sparse attention, the application of sparse activation to the attention mechanism, often face similar challenges.
[…]
@@ 1 Introduction
The machine learning field has experienced a rapid expansion in the scale of Transformer models (Anil et al., 2023; Almazrouei et al., 2023; Dubey et al., 2024; Adler et al., 2024), significantly advancing the state-of-the-art in language understanding and generation. However, this pursuit of larger models is often constrained not by inherent limitations in model quality (Kaplan et al., 2020), but by the soaring computational costs (Sharir et al., 2020; Patterson et al., 2021) associated with the number of parameters. This challenge is compounded by the trend towards models capable of processing increasingly longer input sequences (Reid et al., 2024), where computational demands scale proportionally with input length.
@@
[…]
Activation sparsity has emerged as a prominent technique for mitigating the computational burdens associated with both large model sizes and long input sequences. For large models, activation sparsity reduces computational cost by activating only a small fraction of the model’s parameters for each input. This approach has attracted considerable attention due to the observed lazy neuron phenomenon (Li et al., 2022) where natural activation sparsity occurs in the feed-forward networks (FFNs) of traditional Transformer models like T5 (Raffel et al., 2020) and ViT (Dosovitskiy, 2020), without explicit enforcement (Zhang et al., 2022). Subsequent work has successfully demonstrated wall-time benefits from activation sparsity on multiple hardware platforms, including CPU (Zeng et al., 2023), GPU (Song et al., 2023; Liu et al., 2023), and TPU (Yerram et al., 2024).
Despite the great success, a fundamental challenge arises in the application of sparse activation to the latest and state-of-the-art models. That is, the inherent sparsity, which is pivotal for obtaining efficiency, is largely absent due to the adoption of gated non-ReLU activation functions (Dauphin et al., 2017), widely adopted in, e.g., Mistral (Jiang et al., 2023), Gemma (Gemma Team, 2024), LLAMA (Dubey et al., 2024). This motivates the following question of reactivating sparsity:
@@
To answer this question, prior work has explored reverting to ReLU and its variants (Mirzadeh et al., 2023; Zhang et al., 2024), incorporating top-$k$ thresholding (Yerram et al., 2024; Song et al., 2024a, b), or introducing a low-cost sparsity predictor (Liu et al., 2023; Zeng et al., 2023; Song et al., 2023; Yerram et al., 2024). However, (Challenge #1) these methods often lead to a degradation in model quality (see Figure˜1(a)). Furthermore, (Challenge #2) top-$k$ thresholding based methods not only introduce non-differentiability that compromises model quality, but also involve sorting to obtain maximally activated neurons, which is inefficient on ML accelerators, such as TPUs (Jouppi et al., 2023), with possibly a $10\times$ training slowdown (see Figure˜6). Finally, (Challenge #3) the inclusion of a sparsity predictor often increases training pipeline complexity and incurs additional training costs and parameters.
Sparse attention, the application of activation sparsity to the attention mechanism, presents similar challenges. A common strategy is top-$k$ attention (Gupta et al., 2021), which applies a top-$k$ mask to the attention coefficients. This can be augmented with a low-cost predictor to further enhance efficiency (Ribar et al., 2023; Yang et al., 2024; Lee et al., 2024c). Nevertheless, achieving both high sparsity and accurate prediction without resorting to complex procedures, training slowdown from top-$k$, and compromising quality remains an open problem.
@@
Specifically, the Spark Transformer comprises Spark FFN and Spark Attention, both of which exploit the interpretation of FFNs and attention mechanisms as key-value lookup tables (Geva et al., 2021) to provide a unified framework for sparsity and a low-cost predictor (see Figure˜2). Our low-cost predictor is constructed by repurposing a subset of the dimensions of the query and key vectors to compute an importance score for each key-value pair (see Section˜2). This design addresses Challenge #3 by avoiding the introduction of extra parameters, enabling all model parameters to be trained in a single stage.
Our key technical tool for introducing sparsity is a statistical top-$k$ operator, which is applied to the predicted scores in Spark FFN and Spark Attention to select the activated keys. In particular, statistical top-$k$ is a linear-complexity algorithm for approximate nearest neighbor search, which addresses the issue of high computation cost of standard top-$k$ algorithms (i.e., Challenge #2). As explained in Section˜3, this is achieved by fitting a Gaussian distribution to the activation scores and estimating a threshold that selects the top entries. Moreover, it enjoys continuously differentiability almost everywhere by the usage of a soft-thresholding operator, making the network more amenable to optimization. [^3] Statistical top-k may also be applicable to Mixture-of-Experts, which face the same non-differentiability and inefficiency challenges in their routing mechanism. A recent paper (Wang et al., 2025) also used the soft-thresholding in MoE but still relies on a costly sorting operator. Hence, the linear complexity of statistical top-$k$ could address a key computational bottleneck, especially as MoEs trend toward a larger number of experts (Dai et al., 2024; He, 2024).
Finally, we demonstrate that Spark Transformer alleviates quality loss (i.e., Challenge #1) by performing a full pretraining using the Gemma-2 recipe (Gemma Team, 2024). Comparison with prior work shows that Spark Transformer exhibits a more favorable trade-off between FLOPs reduction and quality measured by pretraining loss (Figure˜1(a)). Further evaluation on standard benchmarks confirms that Spark Transformer closely matches the performance of Gemma-2 (see Figure˜1(b)), despite exhibiting a high degree of sparsity: only 8% of FFN neurons are activated, and each token attends to at most 256 tokens. Leveraging this sparsity, we assess the model’s inference efficiency on CPUs, demonstrating wall-time speedups of $1.86\times$ for prefill and $1.64\times$ for decode (see Figure˜1(c)). Furthermore, an up to $1.4\times$ wall-time speedup is obtained on NVIDIA T4 GPU (see Section˜4). This enhanced efficiency broadens access to high-quality models for users with limited access to high-FLOP hardware, such as high-end GPUs and TPUs.
@@ 2.1 Spark FFN
Each matrix multiplication in Equation˜1 has $2d_{\text{model}}\cdot d_{\text{ff}}$ FLOPs hence overall the computation cost is $4d_{\text{model}}\cdot d_{\text{ff}}$. Previous work shows that when $\sigma()$ is ReLU, the activation map $\sigma({\bm{K}}^{\top}{\bm{q}})$ is very sparse after model training. The sparsity can be used trivially to reduce the computation costs in the calculation of its product with the second layer weight matrix $\bm{V}$} (Li et al., 2022), reducing the overall FLOPs count of FFN to $2d_{\text{model}}\cdot(d_{\text{ff}}+k)$, where $k\ll d_{\text{ff}}$ is the number of nonzero entries in the activation. Note that the sparsity cannot be used to reduce the computation costs associated with $\bm{K}$}, which constitute half of the total FLOPs in FFN.
Figure 2: Architecture of Spark FFN and Spark Attention. (Left) Unified illustration of standard FFN (i.e., Equation˜1) and standard Attention (i.e., Equation˜6). In the case of FFN, ${\bm{q}}\in\mathbb{R}^{d_{\text{model}}}$ is the input, $\bm{K}$} and $\bm{V}$} are the first and second layer weights, respectively, and $\sigma()$ is GELU. In the case of Attention, ${\bm{q}}\in\mathbb{R}^{d_{\text{attn}}}$ is the query, $\bm{K}$} and $\bm{V}$} are key and value matrices, respectively, and $\sigma()$ is softmax. (Right) Unified illustration of Spark FFN (i.e., Equation˜2) and Spark Attention (i.e., Equation˜7). In the case of Spark FFN, $\sigma_{1}()$ is GELU and $\sigma_{2}()$ is identity. In the case of Spark Attention, $\sigma_{1}()$ is softmax and $\sigma_{2}()$ is softplus. In both cases, $\operatorname{{Statistical-Top}}_{k}$ (i.e., Equation˜10) is applied to introduce sparsity, which enables sparse matrix multiplication with ${\bm{K}}_{2}$ and $\bm{V}$} that reduces FLOPs count.
@@
In above, ${\bm{K}}_{1}\subseteq\mathbb{R}^{r\times d_{\text{ff}}}$, ${\bm{K}}_{2}\subseteq\mathbb{R}^{(d_{\text{model}}-r)\times d_{\text{ff}}}$, and ${\bm{V}}\subseteq\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$ are trainable parameters, and the activation $\sigma()$ is taken to be GELU (Hendrycks and Gimpel, 2016) following Gemma. $\operatorname{Top}_{k}$ is introduced for obtaining sparsity, with $k$ being a hyper-parameter specifying the sparsity level. Concretely, $\operatorname{Top}_{k}$ preserves the largest $k$ values in the activation vector, while setting other values to 0. In this paper, we use the efficient statistical top-k presented in Section˜3, which avoids sorting activation values. Finally, the input $\bm{q}$} is split into ${\bm{q}}[:\!r]$ and ${\bm{q}}[r\!:]$, which contain the first $r$ and the rest of the dimensions, respectively, with $r$ being a hyper-parameter. It is introduced so that the term ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ serves as a low-rank predictor of the location of the nonzero entries, which allows us to obtain efficiency benefits in computing ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ and the multiplication with $\bm{V}$}. This is discussed in detail below.
FLOPs per Token. Direct implementation of the Spark-FFN without exploiting sparsity has the same number of FLOPs as the standard FFN in Equation˜1, i.e.,
@@
where the three terms are from multiplication with ${\bm{K}}_{1}$, ${\bm{K}}_{2}$, and $\bm{V}$}, respectively. In Spark-FFN, one may first compute the term ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ as a low-rank predictor. After passing its output through $\operatorname{Top}_{k}$, which selects approximately the $k$ most important entries, followed by the activation function $\sigma()$, we obtain a sparse output. Importantly, after obtaining the sparse output there is no need to perform the full computation of the other two matrix multiplications in Equation˜2, i.e., ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ and the multiplication with $\bm{V}$}. Instead, one can perform a sparse matrix multiplication with a drastically reduced FLOPs count:
@@
which is an increasing function of $r$. In other words, $r$ controls the computation cost. We provide ablation study in Section˜C.4 to show that the best model quality is obtained when $r\approx\frac{d_{\text{model}}}{2}$. In this case, the total FLOP count of Spark FFN is approximately $d_{\text{model}}\cdot d_{\text{ff}}+3\cdot d_{\text{model}}\cdot k$, which is a $4$-times reduction from Equation˜3 when $k$ is small.
Relation to gated activation. Many of the most recent Transformers, including Gemma, use a variant of the standard FFN in Equation˜1 where the activation function is replaced with a gated one:
@@
In above, $\{{\bm{K}}_{1},{\bm{K}}_{2},{\bm{V}}\}\subseteq\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}^{\prime}}$. Note that when compared with the FFN in Equation˜1 for quality studies, $d_{\text{ff}}^{\prime}$ is usually taken to be $2/3\cdot d_{\text{ff}}$ to be iso-parameter count (Shazeer, 2020).
Our Spark FFN in Equation˜2 bears some resemblance to Gated FFN in that both have two linear maps in the first layer and one in the second layer. The difference lies in that 1) Spark FFN adds a $\operatorname{Top}_{k}$ to obtain sparsity, and 2) the input to the first layers of Spark FFN are obtained from splitting the dimensions of the input.